OpenStudy (anonymous):

calculus : find volume of revolution for region bounded by y = sqrt x, y = 0, x = 4, x = 6 , about the line x = 6 (using shell method i think)

OpenStudy (watchmath):

\(\int_4^6 2\pi(6-x)\sqrt{x}\,dx\)

OpenStudy (anonymous):

the book says 0 to 4, must be a mistake

OpenStudy (watchmath):

well if the problem is written exactly what you wrote above the integral should be from x=4 to x=6.

OpenStudy (toxicsugar22):

contrasat can u help me

OpenStudy (toxicsugar22):

please

OpenStudy (toxicsugar22):

4x^2-6y^2+32x-24y+16=0 give the exact coordinated of an four points

OpenStudy (toxicsugar22):

i will give u a medal

OpenStudy (anonymous):

whats the question

OpenStudy (toxicsugar22):

4x^2-6y^2+32x-24y+16=0 give the exact coordinated of an four points

OpenStudy (anonymous):

coordinated is not a word

OpenStudy (anonymous):

of an 4 points? thats not proper english grammar

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

can u help me

OpenStudy (anonymous):

please i need better english

OpenStudy (toxicsugar22):

4x^2-6y^2+32x-24y+16=0 give the exact coordinates of any four points

OpenStudy (toxicsugar22):

on the hyperbola

OpenStudy (anonymous):

ok first change that to a standard hyperbola

OpenStudy (anonymous):

4 ( x^2 - 8x + ? ) - 6 ( y^2 +4 + ? ) + 16 = 0

OpenStudy (toxicsugar22):

ok

OpenStudy (anonymous):

mistake 4 ( x^2 + 8x + ? ) - 6 ( y^2 +4 + ? ) + 16 = 0

OpenStudy (toxicsugar22):

ok

OpenStudy (anonymous):

4 ( x^2 + 8x + (8/2)^2 - (8/2)^2 ) - 6..

OpenStudy (anonymous):

theres an easier way to do this

OpenStudy (anonymous):

just plug in x = 0, and see what you get for y , then x = 1, etc

OpenStudy (anonymous):

4x^2-6y^2+32x-24y+16=0 so if x = 0, we have 0 - 6y^2 + 0 - 24 y + 16 = 0

OpenStudy (toxicsugar22):

ok

OpenStudy (anonymous):

then solve that quadratic

OpenStudy (toxicsugar22):

ok so this is the quadractic -6y^2-24y+16=0

OpenStudy (toxicsugar22):

is that right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok these are not easy coordinates this way

OpenStudy (toxicsugar22):

can i use foil

OpenStudy (anonymous):

you can use quadratic formula, but this is not an easy approach to find 4 coordinates

OpenStudy (toxicsugar22):

can u help me

OpenStudy (anonymous):

YES!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

where is my medal

OpenStudy (toxicsugar22):

i got 92 and 68 but i think i did it wrong can u help me

OpenStudy (anonymous):

where is my medal

OpenStudy (toxicsugar22):

i gave u it

OpenStudy (anonymous):

YAYYYYYYYYYYYYYYYYY!!!!!!

OpenStudy (toxicsugar22):

can u help me

OpenStudy (toxicsugar22):

do the quadractic

OpenStudy (toxicsugar22):

please

OpenStudy (anonymous):

its pretty bad

OpenStudy (toxicsugar22):

what

OpenStudy (anonymous):

i get y = funny bananas

OpenStudy (anonymous):

too much work man

OpenStudy (toxicsugar22):

can u help me

OpenStudy (anonymous):

here, lets do it anothe rway

OpenStudy (anonymous):

YYYYYYYYYYYEEEEESSSSSSSS!!!!!

OpenStudy (toxicsugar22):

can u please help me with the question

OpenStudy (anonymous):

YYYYYYYYYYYYYYYYYYYEEEEEEEEEEEEEEEEEEEEEEEESSSSSSSSSSSSSSSSSSS \[\prime \prime \prime \prime \prime \prime \prime ^{^{^{^{ddd ^{^{^{^{ddd \sqrt[\sqrt[\sqrt[\sqrt[?]{?}]{?}]{?}]{?}}}}}}}}}\]

OpenStudy (toxicsugar22):

are u going to help me or no casue it is rude

OpenStudy (anonymous):

4( x^2 + 8x + (8/2)^2 - (8/2)^2 -

OpenStudy (anonymous):

YYYYYYYYYYYYYYYYYYYYYYYYYYYYYEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEE

OpenStudy (anonymous):

one moment, stand by

OpenStudy (anonymous):

4( x^2 + 8x + (8/2)^2 - (8/2)^2) - 6 ( y^2 + 4y +(4/2)^2 - (4/2)^2 ) + 16 = 0

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

i dont mean to rush u but can u hurry up

OpenStudy (anonymous):

im trying my best

OpenStudy (anonymous):

are you girl or boy

OpenStudy (toxicsugar22):

a boy

OpenStudy (anonymous):

4 ( (x + 4)^2 - 4^2 ) - 16 ( (y+2)^2 - 4 ) + 16 = 0 , agreed?

OpenStudy (anonymous):

i completed the square

OpenStudy (toxicsugar22):

yes

OpenStudy (toxicsugar22):

can u help me now finding the four pointd

OpenStudy (toxicsugar22):

points

OpenStudy (anonymous):

YYYYYYYYYYYYYYYYYYYYEEEEEEEEEEEESSSSSSSSSSSSSSS

OpenStudy (anonymous):

4 ( (x + 4)^2 - 4^2 ) - 16 ( (y+2)^2 - 4 ) + 16 = 0 4 ( x+4)^2 - 4^3 - 16 ( y+2)^2 + 16*4 + 16 = 0

OpenStudy (toxicsugar22):

ok

OpenStudy (anonymous):

4 ( x + 4) ^2 - 16 ( y+2)^2 = 4^3 -16*4 - 16

OpenStudy (anonymous):

4 ( x+4)^2 - 16 (y+2))^2 = -16

OpenStudy (anonymous):

-(x+4)^2 / 4 + (y+2)^2 / 1 = 1

OpenStudy (anonymous):

now solve for y

OpenStudy (anonymous):

(y+2)^2 = ( x+4)^2/4 , square root both sides

OpenStudy (anonymous):

so abs (y+2) = abs(x+4)/2

OpenStudy (anonymous):

+ - (y+2) = + - (x + 4)/2

OpenStudy (toxicsugar22):

so what are the points

OpenStudy (toxicsugar22):

u making me cry

OpenStudy (anonymous):

we're almost there

OpenStudy (toxicsugar22):

what are the points

OpenStudy (toxicsugar22):

please

OpenStudy (anonymous):

i made a mistak,e one sec

OpenStudy (anonymous):

YOURE MAKING ME CRY< lol

OpenStudy (anonymous):

dont cry, just take a deep breath , drink some water

OpenStudy (toxicsugar22):

ok

OpenStudy (anonymous):

ok, i have to go back and find the mistake one sec

OpenStudy (anonymous):

fresh piece of paper

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

i postef a new post

OpenStudy (anonymous):

ok NOOOOOOOOOOOOOOOOOOOOOOOOOOO!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (toxicsugar22):

what

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(y+2)^2 /1^2 - (x+4)^2 / 4 = 1

OpenStudy (anonymous):

(y+2)^2 = 1 + (x+4)^2 /4

OpenStudy (anonymous):

y = + - sqrt ( 1 + (x+4)^2/4)